# Jaime Fernández Fisac and Chen Lv

### v2

@@ Abstract
Dataset. We conduct the experiments on the Waymo Open Motion Dataset (WOMD) (Ettinger et al., 2021), which comprises 486,995 nine-second logged real-world traffic scenarios for training and 44,097 scenarios for validation. The dataset provides trajectories for all agents and corresponding vectorized maps for each scenario. In the Waymo Sim Agents task (Montali et al., 2023), we evaluate our model’s closed-loop simulation performance across 44,920 testing scenarios. In addition, we select 500 scenarios from the WOMD Validation Interactive split for additional experiments and ablation studies.
Model Inputs. For the scene conditions, we consider $A=32$ agents, $M_{l}=256$ map polylines (each containing $M_{p}=30$ waypoints), and $M_{t}=16$ traffic lights in the scene. The flexibility of Transformer blocks allows VBD to be adapted to any number of agents during inference. VBD generates ${T}=80$ steps of future control sequences with step size $0.1s$ based on only the current state and up to 1 s of past trajectories ($T_{h}=11$). Empirically, we observe discarding entire past trajectories and only keeping the current state during inference leads to better closed-loop performance by addressing potential the causal confusion issue in closed-loop testing (Cheng et al., 2024). The effect of history dropout is ablated in Section 8.3.
@@
5 5. VBD as a Closed-loop Traffic Simulator
@@
In this section, we demonstrate the VBD’s capability and generality in generating interactive and realistic traffic scenarios through a closed-loop evaluation using the large-scale Waymo Open Motion Dataset (WOMD) (Ettinger et al., 2021). For each scene, we initialize agent states from WOMD and leverage the Waymax simulator (Gulino et al., 2023) for scenario roll-out. At each time of inference, VBD generates 8-second plans for all agents in the scene via the standard DDPM denoising scheme. All agents replan at 1 Hz using a receding horizon approach. For details on the closed-loop traffic simulation process, refer to Algorithm 2. Furthermore, our qualitative simulation results in Figure 2 highlight VBD’s ability to simulate high-quality traffic scenarios in complex intersections with diverse agent counts, ensuring adherence to maps and traffic controls.
Algorithm 2 Closed-loop traffic simulation with VBD
1:Encoder $\mathcal{E}_{\phi}$, Denoiser $\mathcal{D}_{\theta}$, Initial Scene $\mathbf{c}_{0}$, Denoising Steps $K$, Simulation Steps $T_{s}$, Replan Steps $\Delta t_{r}$.
2:for $t\leftarrow 0,\dots,T_{s}$ do
3:     $t_{r}=t\mod\Delta t_{r}$triangleright Relative time for replanning
4:     if $t_{r}=0$ then $\triangleright$ Check if time to replan
5:         $\hat{\mathbf{c}}_{t}\leftarrow\mathcal{E}_{\phi}(\mathbf{c}_{t})$triangleright Encode scene context
6:         $\tilde{\mathbf{u}}_{K}\sim\mathcal{N}(0,\mathbf{I})$triangleright Sample random noise
7:         for $k\leftarrow K,\dots,1$ do
8:              $\hat{\mathbf{u}}_{0}\leftarrow\mathcal{D}_{\theta}(\tilde{\mathbf{u}}_{k},\hat{\mathbf{c}}_{t},k)$triangleright Denoise
9:              Sample $\tilde{\mathbf{u}}_{k-1}$ using Equation 10
10:         end for
11:     else
12:         Continue using previous control inputs $\tilde{\mathbf{u}}_{0}$
13:     end if
14:     $\mathbf{c}_{t+1}=\text{Step}(\mathbf{c}_{t},\tilde{\mathbf{u}}_{0}^{t_{r}})$triangleright Step simulator with generated actions
15:end for
5.1 5.1 Results on Waymo Sim Agents Challenge
@@
We begin by evaluating the scalability of VBD’s generation quality and interaction modeling capabilities using the Waymo Open Dataset Sim Agents Benchmark (Montali et al., 2023), which comprises 44,920 testing scenes from WOMD. In each scene, we are required to simulate 32 independent closed-loop rollouts of future scenarios for up to 128 agents over an 8-second horizon, conditioning on their trajectories from the preceding 1-second and the corresponding map contexts. Throughout the evaluation, VBD controls up to 64 agents surrounding the labeled self-driving vehicle in each scene, while a constant velocity policy governs any additional agents. By adopting this approach, we significantly reduce computational requirements when sampling millions of scenarios, leveraging the fact that most WOMD scenes contain fewer than 64 dynamic agents, with the remainder being stationary. In practice, we only utilize the current state of all agents by dropping out their past trajectories to overcome the causal confusion. The effectiveness of this strategy will be discussed in Section 8.4.
Evaluation Metrics. We follow the official evaluation metrics of the Waymo Sim Agents benchmark (Montali et al., 2023) encompassing kinematic, interactive, and map-based realism metrics, and a meta realism metric is calculated as a weighted sum of these features. Specifically, kinematic tests measure the likelihood of simulated agents’ linear and angular speed and acceleration compared to the ground truth. The minimum average displacement error (minADE) is also utilized to quantify the positional accuracy relative to the original scenarios.
Table 1: Testing Results on the 2024 Waymo Sim Agents Benchmark
Kinematics $\uparrow$ Interaction $\uparrow$ Map-Based $\uparrow$
Model
Realism
Meta $\uparrow$
Linear
Speed
Linear
Accel
Ang.
Speed
Ang.
Accel
Dist to
Obj
Collision TTC
Dist to
roadedge
Offroad
minADE
[m]$\downarrow$
Logged Oracle - 0.476 0.478 0.578 0.694 0.476 1.000 0.883 0.715 1.000 -
SMART 0.751 0.365 0.406 0.423 0.584 0.377 0.966 0.832 0.659 0.936 1.545
BehaviorGPT 0.747 0.361 0.337 0.481 0.554 0.383 0.954 0.831 0.670 0.935 1.415
GUMP 0.743 0.357 0.411 0.509 0.635 0.371 0.940 0.828 0.669 0.903 1.604
MVTE 0.730 0.351 0.353 0.497 0.600 0.374 0.905 0.831 0.666 0.907 1.677
TrafficBotsV1.5 0.698 0.336 0.350 0.451 0.584 0.360 0.808 0.821 0.642 0.913 1.882
SceneDiffuser 0.703 0.310 0.389 0.459 0.560 0.349 0.917 0.815 0.634 0.833 1.767
VBD (Ours) 0.720 0.359 0.366 0.420 0.522 0.368 0.934 0.815 0.651 0.879 1.474
Table 2: Closed-loop Simulation Results of VBD on Reactive Simulation
Ego planner Model Collision w/ ego [%] Off-road [%] ADE [$m$]
Log-playback Marginal Predictor 10.60 4.49 0.979
VBD Denoiser 4.80 1.43 1.082
IDM-route Marginal Predictor 13.20 5.38 1.070
VBD Denoiser 8.40 2.26 1.107
@@
Results. We evaluate the performance of both diffusion generator and behavior predictor in the Waymo Sim Agents benchmark, and the results are presented in Table 1. We compare the performance of VBD with state-of-the-art autoregressive generative models, including BehaviorGPT (Zhou et al., 2024), SMART (Wu et al., 2024), and GUMP (Hu et al., 2024), behavior cloning models, such as MVTE (Wang et al., 2023b) and TrafficBotsV1.5 (Zhang et al., 2024), and concurrent diffusion-based model SceneDiffuser (Jiang et al., 2024). We demonstrate that the simulation performance of the VBD model is comparable to state-of-the-art autoregressive generative models and outperforms other motion prediction models. In addition, we observe better metrics across almost all categories compared to SceneDiffuser (Jiang et al., 2024) baseline from Waymo. Notably, our VBD model achieves this with fewer parameters than these GPT-style models, achieving a balance between performance and computational efficiency. We present a selection of qualitative simulation results in Figure 2, showcasing the model’s ability to generate interactive and realistic traffic scenarios. The results in the Waymo Sim Agents task suggest that our VBD model is reliable at generating realistic agent interactions through its joint multi-agent diffusion policy. Additional simulation videos can be found on the Project Website.
5.2 5.2 Reactive Simulation with AV Planner
A critical aspect of traffic simulation is ensuring realistic agent responses to the ego vehicle’s maneuvers, regardless of the ego’s underlying policy. This experiment aims to quantify the impact of scene-centric representation in VBD on simulation realism. We conduct reactive simulation tests, where the ego vehicle is controlled by an external planner while surrounding agents are controlled by our models. The evaluation uses a subset of 500 WOMD interactive validation scenarios, previously selected by Zhang et al. (2023a). Each simulation selects up to 32 agents from the original scene, excluding any additional agents, and simulations run for 8 seconds with a 1 Hz replanning frequency. For a fair comparison, we use the co-trained marginal behavior predictor as a baseline. This predictor shares the same environment context encoding and similar architecture as the denoiser. In the baseline policy, we sample maximum likelihood trajectories for each agent and replan at 1 Hz.
Evaluation Metrics. We employ several metrics from the Waymax simulator (Gulino et al., 2023) to evaluate the performance of our model. Specifically, we track three key indicators:
  • Off-road incidents: The percentage of agents that deviate from designated drivable areas within the map.
  • Collisions with ego: The frequency of collisions between agents and the ego vehicle, reflecting the responsiveness of the behavior simulation engine.
  • Log divergence: A measure of the discrepancy between simulated and recorded behaviors, calculated using the average displacement error (ADE).
Results. We assessed the reactivity of our model to the ego vehicle’s actions under two distinct control scenarios: a log-playback from WOMD and an IDM-route planner. According to our findings, summarized in Table 2, both the marginal predictor and VBD denoiser demonstrated comparable levels of log divergence (ADE), indicating their ability to mimic realistic trajectories. However, the diffusion joint policy showed notable superiority over the marginal behavior prediction approach in terms of reactivity, yielding fewer collisions with the ego vehicle. Interestingly, when utilizing the IDM-route planner – known for its deviations from human-like driving patterns and actual trajectories – the performance suffered. Conversely, employing a log-playback planner, which closely resembles the training data distribution, enabled our VBD model to exhibit highly reactive behaviors in response to the ego vehicle’s actions.
6 6. VBD as a Controllable Scenario Generator
@@
A key strength of the VBD model lies in its capacity to generate customizable traffic scenarios without requiring retraining. This flexibility stems from the guided generation, allowing the model to adapt behaviors based on user-specified objectives, such as avoiding collisions or creating adversarial scenarios. By integrating these objectives into the denoising process, the VBD model provides a versatile and potent tool for simulating diverse traffic scenarios on demand. As outlined in Algorithm 3, users can define cost objectives $\mathcal{J}_{y}$, enabling compositional optimization via the classifier guidance method (Dhariwal and Nichol, 2021). This involves iteratively refining the noisy $\tilde{\mathbf{u}}$ using both denoiser outputs and gradients of the form $\nabla{\tilde{\mu}_{k}}\mathcal{J}_{y}(\mathcal{D}\theta(\tilde{\mu}_{k}))$. Throughout this process, the denoiser ensures scenario realism, while the user’s objective encourages the model to explore desired modes. Subsequent sections will demonstrate the application of VBD’s capabilities across various simulation tasks. Specifically, our guided generation tasks utilize $N_{g}=5$ gradient steps and a fixed scaling parameter $\lambda=0.1$ during all denoising operations.
@@
2:$\tilde{\mathbf{u}}_{K}\sim\mathcal{N}(0,\mathbf{I})$triangleright Sample initial trajectory
3:for $k\leftarrow K$ to $1$ do
4:     $\hat{\mathbf{u}}\leftarrow\mathcal{D}_{\theta}(\tilde{\mathbf{u}}_{k},k,\mathbf{c})$triangleright Predict denoised control sequence
5:     ${\tilde{\mu}}_{k}\leftarrow\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\tilde{\mathbf{u}}_{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\hat{\mathbf{u}}$triangleright Calculate unguided posterior $\tilde{\mu}_{k}$
@@
7:         ${\tilde{\mu}}_{k}\leftarrow{\tilde{\mu}}_{k}+\lambda\sigma_{k}\nabla_{\tilde{\mu}_{k}}\mathcal{J}_{y}(\mathcal{D}_{\theta}(\tilde{\mu}_{k}))$triangleright Guidance gradient step
@@
9:     $\tilde{\mathbf{u}}_{k-1}\sim\mathcal{N}(\tilde{\mu}_{k},\sigma_{k}^{2}\mathbf{I})$triangleright Sample previous-step noised control sequence
@@
Setup. We showcase the controllable scenario generation capabilities of our model in fulfilling various simulation tasks. To test this, we leverage the same subset of 500 WOMD interactive validation scenarios employed in Section 5.2. Within each scene, we select the 32 agents closest to the labeled ego agent for evaluation, excluding all others. Our simulations span an $8$-second horizon with a replanning frequency of $1$ Hz, utilizing metrics from the Waymax simulator including off-road incidents, collisions, wrong-way occurrences, kinematic infeasibilities, and log divergences. Furthermore, we analyze all quantitative results across three distinct random seeds, averaging outcomes to assess the generality and consistency of VBD.
Evaluation Metrics. To evaluate the performance of our proposed model, we provide both qualitative and quantitative results. Additional qualitative results including more illustrative video examples can be found on the Project Website. Furthermore, We employ with following metrics to thoroughly test our proposed method quantitatively:
  • Off-road: A binary metric indicating whether a vehicle drives off the road, determined by its position relative to oriented road graph points. A vehicle is considered off-road if it strays from the right side of an oriented road edge. The off-road rate is calculated as the percentage of vehicles transitioning from on-road to off-road across all simulated agents and generated scenarios.
  • Collision: A binary metric identifying collisions between agents. Collisions are detected when the 2D bounding boxes of two objects overlap at the same time step. The collision rate is calculated as the percentage of agents involved in collisions across all simulated agents and generated scenarios.
  • Wrong-way: A binary metric measuring whether a vehicle deviates from its intended driving direction. A wrong-way movement is flagged if the vehicle’s heading angle deviates more than 90 degrees from its closest lane direction for over 1 second. The wrong-way rate is calculated as the percentage of vehicles entering the wrong way across all simulated vehicles and generated scenarios.
  • Kinematic infeasibility (Kin.): A binary metric assessing whether a vehicle’s transition steps are kinematically plausible. The limits for acceleration magnitude (6 $m/s^{2}$) and steering curvature magnitude (0.3 $m^{-1}$) are set empirically. The kinematic infeasibility rate is calculated as the percentage of vehicles with infeasible trajectories across all simulated vehicles and generated scenarios.
  • Log divergence (ADE and FDE): These metrics quantify the deviation from ground truth behavior using displacement error. Average Displacement Error (ADE) is the L2 distance between an agent’s simulated and ground-truth positions at each time step. Final Displacement Error (FDE) is the L2 distance between an agent’s simulated and ground-truth positions at $t=8$ seconds. ADE and FDE are reported as averages across all time steps and valid agents in all generated scenarios.
  • Minimum Log divergence (minADE and minFDE): These metrics quantify the minimum deviation from ground truth behavior using displacement error among all generated scenarios for each scene. MinADE and minFDE are reported as averages across all time steps and valid agents for all scenes.
Figure 4: VBD produces scene-consistent traffic scenarios when target agents are conditioned on specific behavior priors. Top: Nominal VBD rollout without guidance generates a scene-consistent scenario, where Vehicle 5 (in purple) waits at the stop sign and then precedes. Bottom: Using goal-guided diffusion to minimize Vehicle 5’s final position w.r.t to a given goal, we enforce Vehicle 5 to run the stop sign. VBD model generates a scene-consistent scenario with Vehicle 1 yielding to Vehicle 5.
Table 3: Simulation Results of Scenario Editing with Marginal Predictor and Diffusion Policy over Three Different Seeds
@@
Marginal Prediction 5.61$\pm$0.27 6.19$\pm$0.05 0.86$\pm$0.09 0.31$\pm$0.02 1.113$\pm$0.012
Diffusion 2.47$\pm$0.09 1.21$\pm$0.14 0.57$\pm$0.02 0.24$\pm$0.01 1.010$\pm$0.007
Post-Optimization via Diffusion 2.23$\pm$0.15 1.26$\pm$0.10 0.52$\pm$0.11 0.32$\pm$0.01 0.974$\pm$0.005
6.1 6.1 Composition with Behavior Priors
Sampling diverse outputs from a conditional diffusion model is challenging, especially when the denoising strongly relies on the context information (Sadat et al., 2023). On the other hand, the behavior predictor in VBD captures the multi-modal trajectories of individual agents but will result in scene inconsistency if marginal trajectories are naively combined. This is because the predictor alone cannot ensure the coherence necessary for realistic multi-agent scenario generation. The denosier in VBD can used as an effective scenario optimizer and produce diverse and scene-consistent scenarios by first selecting goal positions for some target agents and generating joint multi-agent trajectories matching individual goals using guided sampling. Consider a scenario where we need to specify desired behaviors for $A^{t}$ agents, we can heuristically determine the target behaviors or goals for each target agent, represented by $\{\mathbf{g}^{i}\}_{i=1:A^{t}}$. The guidance function based on behavior priors is:

### v3

@@ V-A System Dynamics
The unicycle dynamics used for mapping between actions and states may introduce small discrepancies between reconstructed and ground-truth trajectories. However, our empirical analysis shows that this effect is minor in practice. Across 40K agent trajectories from the Waymo Motion dataset, the average displacement error between 8-second reconstructed and ground-truth trajectories is 0.221 m, with a final displacement error of 0.511 m. These results show that the inverse dynamics model yields sufficiently accurate state-action reconstructions. In our framework, inverse dynamics is used only to obtain an approximate control-space representation, while supervision is applied on state trajectories after forward rollout through the differentiable forward dynamics function. This rollout-based supervision mitigates potential mismatches, ensuring consistent trajectory generation.
V-B Dataset
We conduct the experiments on the Waymo Open Motion Dataset (WOMD) [14], which comprises 486,995 nine-second logged real-world traffic scenarios for training and 44,097 scenarios for validation. The dataset provides trajectories for all agents and corresponding vectorized maps for each scenario. In the Waymo Sim Agents task [36], we evaluate our model’s closed-loop simulation performance across 44,920 testing scenarios. In addition, we select 500 scenarios from the WOMD Validation Interactive split for additional experiments and ablation studies.
We choose only the WOMD instead of nuScenes [2] and Argoverse [4] for the following reasons. First, nuScenes and Argoverse are primarily designed for perception and short-term prediction benchmarks, which are insufficient for extended, interactive multi-agent simulation. Second, neither provides a standardized closed-loop evaluation framework that can simulate and replay the behavior of all traffic participants over the full duration of each scenario. In contrast, we use the Waymax environment [17] built on WOMD, which supports detailed, longer-horizon (8 seconds) closed-loop simulation of every agent in the scene, allowing us to rigorously assess multi-agent interaction dynamics.
V-C Implementation Details
Model Inputs. For the scene conditions, we consider $A=64$ agents, $M_{l}=256$ map polylines (each containing $M_{p}=30$ waypoints), and $M_{t}=16$ traffic lights in the scene. The flexibility of Transformer blocks allows VBD to be adapted to any number of agents during inference. VBD generates T=80{T}=80 steps of future control sequences with step size $0.1s$ based on only the current state and up to 1 s of past trajectories ($T_{h}=11$). Empirically, we observe that discarding entire past trajectories and only keeping the current state during inference leads to better closed-loop performance by addressing the causal confusion issue in closed-loop testing [8].
@@ V-C Implementation Details
VI Closed-loop Traffic Simulation
@@
We demonstrate the VBD’s capability and generality in generating interactive and realistic traffic scenarios through a closed-loop evaluation. For each scene, we initialize agent states from WOMD and leverage the Waymax simulator [17] for scenario roll-out. At each time of inference, VBD generates 8-second trajectories for all agents in the scene via the standard DDPM denoising scheme. All agents replan at 1 Hz using a receding horizon approach.
@@
We evaluate the scalability of VBD’s generation quality and interaction modeling capabilities using the Waymo Sim Agents Benchmark. In each scene, we are required to simulate 32 independent closed-loop rollouts of future scenarios for up to 128 agents over an 8-second horizon, conditioning on their trajectories from the preceding 1-second and the corresponding map contexts. Throughout the evaluation, VBD controls up to 64 agents surrounding the labeled self-driving vehicle in each scene, while a constant velocity policy controls other agents. By adopting this approach, we significantly reduce computational requirements when sampling millions of scenarios, leveraging the fact that most WOMD scenes contain fewer than 64 dynamic agents, with the remainder being stationary. We only utilize the current state of all agents by dropping out their past trajectories to overcome the causal confusion. We follow the official evaluation metrics of the Waymo Sim Agents benchmark [36], encompassing kinematic, interactive, and map-based metrics, and a meta realism metric is calculated as a weighted sum of these features.
TABLE I: Testing Results on the 2024 Waymo Sim Agents Benchmark
Agent Policy
Realism Meta
Metric ($\uparrow$)
Kinematic
Metric ($\uparrow$)
Interactive
Metric ($\uparrow$)
Map-based
Metric ($\uparrow$)
SMART 0.759 0.476 0.804 0.767
BehaviorGPT 0.747 0.433 0.800 0.764
GUMP 0.743 0.478 0.789 0.740
MVTE 0.730 0.450 0.771 0.743
SceneDiffuser 0.703 0.429 0.749 0.767
TrafficBotsV1.5 0.699 0.430 0.711 0.744
VBD (Ours) 0.720 0.417 0.782 0.721
@@
In Table I, we compare the performance of VBD with state-of-the-art generative models, including BehaviorGPT [66], SMART [57], and GUMP [21], behavior cloning models, such as MVTE [55] and TrafficBotsV1.5 [63], and concurrent diffusion-based model SceneDiffuser [28]. We demonstrate that the simulation performance of the VBD model is comparable to state-of-the-art generative models and outperforms other motion prediction models. In addition, we observe better simulation realism metrics compared to the SceneDiffuser [28] baseline. Notably, our VBD model achieves this with fewer parameters than autoregressive generation models, achieving a balance between performance and computational efficiency. We present a selection of qualitative simulation results in Fig. 3, showcasing the model’s ability to generate interactive and realistic traffic scenarios. The results in the Waymo Sim Agents benchmark suggest that our VBD model is reliable at generating realistic agent interactions through its joint multi-agent diffusion policy. Closed-loop simulation videos can be found on the Project Website.
While the proposed VBD model achieves strong results on the Sim Agent benchmark, a performance gap remains compared to autoregressive models such as SMART and BehaviorGPT. We believe this discrepancy largely comes from architectural and training factors. In particular, the map representation and encoder design in VBD can be further improved to better capture spatial dependencies between the road map and agents’ future trajectories, and increasing the number of denoising steps yields limited benefit under the realism metric. Despite these factors, the VBD model offers distinct advantages, including joint multi-agent consistency and flexible inference-time controllability, enabling user-guided behavior generation such as adversarial maneuvers.
VII Controllable Scenario Generation
@@
A key strength of the VBD model lies in its capacity to generate customizable traffic scenarios without requiring retraining. This flexibility stems from the guided generation, allowing the model to adapt behaviors based on user-specified objectives, such as avoiding collisions or creating adversarial scenarios. By integrating these objectives into the denoising process, the VBD model provides a versatile and potent tool for simulating diverse traffic scenarios on demand. As outlined in Algorithm 1, users can define cost objectives $\mathcal{J}_{y}$, enabling compositional optimization via the classifier guidance method [11]. This involves iteratively refining the noisy $\tilde{\mathbf{u}}$ using both denoiser outputs and gradients of the form $\nabla{\tilde{\mu}_{k}}\mathcal{J}_{y}(\mathcal{D}\theta(\tilde{\mu}_{k}))$. Throughout this process, the denoiser ensures scenario realism, while the user’s objective encourages the model to explore desired modes. Subsequent sections will demonstrate the application of VBD’s capabilities across various simulation tasks. Specifically, our guided generation tasks utilize $N_{g}=5$ gradient steps and a fixed scaling parameter $\lambda=0.1$ during all denoising operations.
@@
[…]
3:for $k\leftarrow K$ to 11 do
4:  $\hat{\mathbf{u}}\leftarrow\mathcal{D}_{\theta}(\tilde{\mathbf{u}}_{k},k,\mathbf{c})$ $\triangleright$ Predict denoised control sequence
5:  $^{\tilde{\mu}}_{k}\leftarrow\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\tilde{\mathbf{u}}_{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\hat{\mathbf{u}}$ $\triangleright$ Calculate unguided posterior $\tilde{\mu}_{k}$
@@
7:   ${\tilde{\mu}}_{k}\leftarrow{\tilde{\mu}}_{k}+\lambda\sigma_{k}\nabla_{\tilde{\mu}_{k}}\mathcal{J}_{y}(\mathcal{D}_{\theta}(\tilde{\mu}_{k}))$ $\triangleright$ Guidance gradient step
@@
9:  $\tilde{\mathbf{u}}_{k-1}\sim\mathcal{N}(\tilde{\mu}_{k},\sigma_{k}^{2}\mathbf{I})$ $\triangleright$ Sample previous-step noised control sequence
@@
Figure 4: VBD produces scene-consistent traffic scenarios when target agents are conditioned on specific behavior priors. Top: Nominal VBD rollout without guidance generates a scene-consistent scenario, where Vehicle 5 (in purple) waits at the stop sign and then proceeds. Bottom: Using goal-guided diffusion to minimize Vehicle 5’s final position w.r.t to a given goal, we enforce Vehicle 5 to run the stop sign. VBD model generates a scene-consistent scenario with Vehicle 1 yielding to Vehicle 5.
TABLE II: Simulation Results of Scenario Editing with Marginal Predictor and Diffusion Policy over Three Random Seeds
@@
[…]
We showcase the controllable scenario generation capabilities of our model in fulfilling various simulation tasks. To test this, we leverage the same subset of 500 WOMD interactive validation scenarios. Within each scene, we select the 32 agents closest to the labeled ego agent for evaluation, excluding all others. Our simulations span an $8$-second horizon with a replanning frequency of $1$ Hz, utilizing metrics from the Waymax simulator, including off-road incidents, collisions, wrong-way occurrences, kinematic infeasibilities, and log divergences. Furthermore, we analyze all quantitative results across three distinct random seeds, averaging outcomes to assess the generality and consistency.
We employ the following metrics to thoroughly test our proposed method quantitatively:
  • Off-road: A binary metric indicating whether a vehicle drives off the road, determined by its position relative to oriented road graph points.
  • Collision: A binary metric identifying if collisions happen between agents.
  • Wrong-way: A binary metric measuring whether a vehicle deviates from its intended driving direction.
  • Kinematic infeasibility (Kin.): A binary metric assessing whether a vehicle’s movement is kinematically plausible.
  • Log divergence (ADE and FDE): These metrics quantify the deviation from ground truth behavior using displacement error. Average Displacement Error (ADE) is the L2 distance between an agent’s simulated and ground-truth positions at each time step. Final Displacement Error (FDE) is the L2 distance error at $t=8$ seconds.
  • Minimum Log divergence (minADE and minFDE): These metrics quantify the minimum deviation from ground truth behavior using displacement error among all generated scenarios for each scene.
VII-A Composition with Behavior Priors
Sampling diverse outputs from a conditional diffusion model is challenging, especially when denoising strongly relies on the context information [43]. On the other hand, the behavior predictor in VBD captures the multi-modal trajectories of individual agents but will result in scene inconsistency if marginal trajectories are naively combined. This is because the predictor alone cannot ensure the coherence necessary for realistic multi-agent scenario generation. The denoiser in VBD can be used as an effective scenario optimizer and produce diverse and scene-consistent scenarios by first selecting goal positions for some target agents and generating joint multi-agent trajectories matching individual goals using guided sampling. Consider a scenario where we need to specify desired behaviors for $A^{t}$ agents, we can heuristically determine the target behaviors or goals for each target agent, represented by $\{\mathbf{g}^{i}\}_{i=1:A^{t}}$. The guidance function with behavior priors is: